# Robust Unsupervised Domain Adaptation for 3D Point Cloud Segmentation Under Source Adversarial Attacks

### v1

@@ III The AdvSynLiDAR Dataset
Generally, point clouds in the source domain are susceptible to subtle perturbations and deformations, which can be caused by sensor noise, environmental variations, and preprocessing artifacts [23, 24]. This can significantly degrade model performance and increase distribution discrepancies across domains, leading to inferior 3D unsupervised cross-domain adaptation (UDA) performance for point cloud semantic segmentation. Despite the significance of this challenge, to the best of our knowledge, there is currently a lack of standardized methods for evaluating the impact of such perturbations during adaptation. To address these issues, we introduce the AdvSynLiDAR dataset, built upon the structured SynLiDAR dataset [3]. Specifically, our dataset focuses on understanding the impact of attacks on the source domain. To simulate such an attack, we assume that we are able to access the point cloud $\mathbf{x}_{\mathcal{S}}$ and labels $\mathbf{y}_{\mathcal{S}}$ in the source domain. To obtain a point-wise classification and confidence output of point cloud, the pre-trained model $\theta$ provides the probability distribution of the $i$-th point in $\mathbf{x}_{\mathcal{S}}$, which is denoted as $\Phi_{\theta}(\mathbf{x}_{\mathcal{S}})^{(i)}=p(\cdot|\mathbf{x}_{\mathcal{S}})^{(i)}\in\mathbb{R}^{|c|}$.
@@ III-A Imperceptible Perturbation Points
In real-world scenarios, adversarial perturbations are usually subtle and spatially dependent, making them difficult to be detected. To mimic realistic attacks, we propose a perceptually-aware perturbation strategy. This strategy applies more significant perturbations to distant regions while minimizing changes in dense, perceptually sensitive areas, enabling a more precise evaluation of UDA robustness under an adversarial source domain. Motivated by the observation that central regions near the viewpoint exhibit higher point density and perceptual sensitivity, we introduce a dynamic adjustment factor $\gamma^{(i)}$ to regulate the magnitude of perturbations based on their proximity to the viewpoint center $C$. For each point $\mathbf{x}_{\mathcal{S}}^{(i)}$ in the point cloud, the distance $d_{i}$ from the viewpoint center is calculated by $d(\mathbf{x}_{\mathcal{S}}^{(i)})=\|\mathbf{x}_{\mathcal{S}}^{(i)}-C\|$, and the dynamic factor $\gamma^{(i)}$ is denoted as:
$\gamma^{(i)}=\frac{d(\mathbf{x}_{\mathcal{S}}^{(i)})}{d_{\max}},$ (1)
where $d_{\max}$ denotes the maximum distance to the viewpoint. The adjusted PGD attack is denoted as:
$\mathbf{x}_{i+1}^{\mathrm{adv}}$ $=\mathrm{clip}_{\mathbf{x},\epsilon}\Big{(}\mathbf{x}_{i}^{\mathrm{adv}}$ (2)
$+\gamma^{(i)}\odot\delta\cdot\mathrm{sign}(\nabla_{\mathbf{x}_{i}^{\mathrm{adv}}}L(f(\mathbf{x}_{i}^{\mathrm{adv}};\theta),\mathbf{y}_{i}))\Big{)},$
where $\theta$ is the parameters of the target model, $\delta$ denotes the perturbation intensities of the original point cloud. Term $\nabla_{\mathbf{x}_{i}^{\mathrm{adv}}}L$ is the gradient value of the adversarial loss function concerning the adversarial point cloud in the $i$-th iteration, and $y_{i}$ corresponds to the ground truth label of the original point. The $\text{clip}(\cdot)$ function constrains the perturbation range of the adversarial point cloud within $[0,\epsilon]$ relative to the original data. As shown in Fig. 2, by dynamically adjusting the magnitude of perturbations, we generate adversarial point clouds that effectively maintain the potency of attack while remaining less detectable.
@@ III-B Noisy Semantic Labels
Manual annotations in 3D point clouds often contain label noise, especially for occluded, distant, or complex objects, where annotation errors occur more frequently [25]. While standard UDA approaches assume that source domain annotations are reliable, real-world scenarios often involve systematic annotation noise, where annotators or automated labeling systems confidently assign incorrect labels. To systematically evaluate the impact of such errors, we introduce high-confidence mislabeling into the source domain data. Specifically, we formalize high-confidence mislabeling by extracting prediction confidence from the model’s logit output: given a source domain point-cloud $\mathbf{x}_{\mathcal{S}}$, and pre-trained model $\theta_{\mathcal{S}}$, the adversarial labels are generated via:
$\tilde{\mathbf{y}}^{(i)}=\underset{\mathbf{y}\in\mathbf{Y}_{\mathcal{S}}\setminus\{\mathbf{y}^{(i)}\}}{\arg\max}\,\{p(\mathbf{y}\mid\mathbf{x}_{\mathcal{S}}^{(i)})\mid p(\mathbf{y}\mid\mathbf{x}_{\mathcal{S}}^{(i)})>\tau\},$ (3)
where $\tau$ denotes the confidence threshold, and $p(\mathbf{y}|\mathbf{x}_{\mathcal{S}}^{(i)};\theta_{\mathcal{S}})$ denotes the predicted class probabilities. This approach generates the highest-confidence incorrect class as the adversarial target, mimicking real-world scenarios with noisy labels.
@@ IV-B1 Robust Long-Tail Loss (RLT Loss)
During the pre-training phase, in order to balance the advantages of KPS loss [5] in handling long-tail classes with the strengths of SoftDICELoss in managing head classes, we introduce the Robust Long-Tail loss (RLT loss). SoftDiceLoss is widely used in segmentation tasks to measure the overlap between predictions and ground truth labels. It is defined as:
$\mathcal{L}_{\text{SD }}=1-\frac{2\sum_{i=1}^{N}\mathbf{p}_{\mathcal{S}}^{(i)}\mathbf{y}_{\mathcal{S}}^{(i)}}{\sum_{i=1}^{N}\left(\mathbf{p}_{\mathcal{S}}^{(i)}\right)^{2}+\sum_{i=1}^{N}\left(\mathbf{y}_{\mathcal{S}}^{(i)}\right)^{2}}.$ (4)
This formula calculates the Dice similarity [26] coefficient between the predicted results $\mathbf{p}_{\mathcal{S}}^{(i)}$ and the ground truth labels $\mathbf{y}_{\mathcal{S}}^{(i)}$, where N is the number of samples. Initially designed for 2D classification tasks, KPS loss aims to enhance the separability of features by adjusting the boundaries between key points and non-key points in the feature space. To apply KPS loss to higher-dimensional 3D segmentation tasks, we make several improvements to the original KPS loss:
Geometric Importance in 3D Key Point Definition. We introduce geometric importance as the criteria for defining key points in 3D point clouds, allowing for better identification of critical points under attack conditions. Specifically, we define the geometric importance of a key point using the following formula:
$\text{Importance}\left(\mathbf{x_{i}}\right)=\frac{\sum_{j=1}^{k}\text{Distance}\left(\mathbf{x_{i}},\mathbf{x_{j}}\right)}{k},$ (5)
where $\mathbf{x_{i}}$ is the current point, $\mathbf{x_{j}}$ are its $k$ nearest neighbors, and $\text{Distance}\left(\mathbf{x_{i}},\mathbf{x_{j}}\right)$ calculates the Euclidean distance between point $\mathbf{x_{i}}$ and $\mathbf{x_{j}}$. This enhances the discriminative ability of KPS loss for long-tail classes.
@@
Dynamic Boundary Adjustment for Class Imbalance. In standard classification tasks, all classes share the same decision boundary, which may not be ideal for long-tail classes due to their sparse distribution and low sample count. However, 3D point clouds often exhibit severe class imbalance, especially for long-tail classes as depicted in Fig. 1. To address this issue, we introduce a dynamic boundary adjustment mechanism that adjusts each class’s boundary based on sample density and classification difficulty:
$m_{y_{i}}=\alpha\cdot\log\left(\frac{CS_{\mathbf{y}_{i}}}{CS_{max}}\right),$ (6)
where $\alpha$ is a scaling factor, ${CS_{\mathbf{y}_{i}}}$ is the sample count for class $\mathbf{y}_{i}$, and ${CS_{max}}$ is the sample count of the largest class. In this way, the smaller the class size (i.e., long-tail classes), the larger the boundary adjustment term $m_{\mathbf{y}_{i}}$, giving these classes more decision space during classification.
In summary, the KPS loss in 3D tasks is defined as:
$\mathcal{L}_{\text{KPS}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{e^{s\cdot\left(\cos\left(\theta_{\mathbf{y}_{i}}\right)-m_{\mathbf{y}_{i}}\right)}}{e^{s\cdot\left(\cos\left(\theta_{\mathbf{y}_{i}}\right)-m_{\mathbf{y}_{i}}\right)}+\sum_{j\neq\mathbf{y}_{i}}e^{s\cdot\cos\left(\theta_{j}\right)}},$ (7)
where $s$ is a scaling factor, $\theta$ is the angle between the feature vector and the class anchor vector, and $m_{\mathbf{y}_{i}}$ is the boundary for class $\mathbf{y}_{i}$. This loss function encourages more significant boundaries for key points, thereby promoting the separability of long-tail class features.
@@ IV-B2 Probability Decoder Branch
Probability Decoder. Inspired by VRCNet [6], we introduce a probability decoder to mitigate latent discrepancies in source inputs dynamically. As illustrated in Fig. 3, to simulate realistic noise, we generate perturbed inputs $\mathbf{x}_{g}$ by adding Gaussian noise and random augmentations to the clean source coordinates $\mathbf{x}_{\mathcal{S}}$, i.e., $\mathbf{x}_{g}=r(\mathbf{x}_{\mathcal{S}}+\mathcal{N}(0,\sigma^{2}))$. The decoder branch reconstructs the clean point cloud in three stages: reconstruction, completion, and enhancement, which progressively restore global and local geometric details. Specifically, in the reconstruction stage, VRCNet encodes the input features $\mathbf{x}_{\mathcal{S}}$ into global features $\mathbf{z}_{g}$ and the latent distribution $q_{\phi}(\mathbf{z}_{g}|\mathbf{x}_{\mathcal{S}})$. It then reconstructs a coarse point cloud $\mathbf{x}_{r}^{\prime}$ using the decoding distribution $p_{\theta}^{r}(\mathbf{x}_{\mathcal{S}}|\mathbf{z}_{g})$. The training loss of reconstruction is denoted as:
$\mathcal{L}_{rec}=$ $-\lambda\mathbf{KL}\left[q_{\phi}(\mathbf{z_{g}}|\mathbf{x}_{\mathcal{S}})\|p(\mathbf{z_{g}})\right]$ (9)
$+\mathbb{E}_{p_{data}(\mathbf{x}_{\mathcal{S}})}\mathbb{E}_{q_{\phi}(\mathbf{z_{g}}|\mathbf{x}_{\mathcal{S}})}\Big{[}\log p_{\theta}^{r}(\mathbf{x}_{\mathcal{S}}|\mathbf{z_{g}})\Big{]},$
where $\mathbf{KL}$ represents the KL divergence loss, and $\mathbb{E}$ denotes the expected value based on the probability distribution. $p_{data}(\mathbf{x}_{\mathcal{S}})$ refers to the coordinate distribution of the clean source, and $p(\mathbf{z}_{g})=\mathcal{N}(\mathbf{0},\mathbf{I})$ denotes the predefined prior distribution, which is a Gaussian distribution. Additionally, the parameter $\lambda$ serves as a weighting factor. During the completion process, the input point cloud $\mathbf{x}_{c}^{\prime}$ is reconstructed using the global feature $\mathbf{z}_{g}$ and the latent variable distribution $p_{\psi}(\mathbf{z}_{g}|\mathbf{x}_{g})$ derived from the perturbed and augmented input point cloud $\mathbf{x}_{g}$. During training, the posterior distribution $p_{\psi}(\mathbf{z}_{g}|\mathbf{x}_{g})$ is regularized to approach the prior distribution $q_{\phi}(\mathbf{z}_{g}|\mathbf{x}_{\mathcal{S}})$. The loss function is thus:
$\mathcal{L}_{com}=$ $-\lambda\mathbf{KL}\left[q_{\phi}(\mathbf{z_{g}}|\mathbf{x}_{\mathcal{S}})\|p_{\psi}(\mathbf{z_{g}}|\mathbf{x_{g}})\right]$ (10)
$+\mathbb{E}_{p_{data}(\mathbf{x_{g}})}\mathbb{E}_{p_{\psi}(\mathbf{z_{g}}|\mathbf{x_{g}})}\big{[}\log p_{\theta}^{c}(\mathbf{x}_{\mathcal{S}}|\mathbf{z_{g}})\big{]},$
in which the paths for completion and reconstruction share identical encoder and decoder weights. In the enhancement stage, a self-attention module and up-sampling layers are added to capture local features at different scales and restore the point cloud to finer granularity [6], denoted $\mathbf{x}_{f}^{\prime}$. To facilitate stable point cloud generation, we replace the original log-likelihood objective $log~{}p(x|z)$ with the Chamfer Distance $\mathcal{L}_{CD}$ [28] which is more suitable for optimizing unordered point sets. The overall training objective of the decoder integrates all three stages:
$\mathcal{L}_{d}=$ $\lambda_{rec}\mathcal{L}_{rec}+\lambda_{com}\mathcal{L}_{com}+\lambda_{fine}\mathcal{L}_{fine}$ (11)
$=$ $\lambda_{rec}\big{[}\mathcal{L}_{\mathbf{KL}}(q_{\phi}(\mathbf{z_{g}}|\mathbf{x}_{\mathcal{S}}),\mathcal{N}(\mathbf{0},\mathbf{I}))+\mathcal{L}_{\mathbf{CD}}(\mathbf{x}_{r}^{\prime},\mathbf{x}_{\mathcal{S}})\big{]}$
$+\lambda_{com}\big{[}\mathcal{L}_{\mathbf{KL}}(p_{\psi}(\mathbf{z_{g}}|\mathbf{x_{g}}),q_{\phi}(\mathbf{z_{g}}|\mathbf{x}_{\mathcal{S}}))+\mathcal{L}_{\mathbf{CD}}(\mathbf{x_{c}^{\prime}},\mathbf{x}_{\mathcal{S}})\big{]}$
$+\lambda_{fine}\mathcal{L}_{\mathbf{CD}}(\mathbf{x}_{f}^{\prime},\mathbf{x}_{\mathcal{S}}).$
This enables adaptive restoration of perturbed point clouds while preserving essential structural information, mitigating the risk of introducing additional noise or erroneous labels.
@@
High-Confidence Nearest Pseudo Update. Motivated by existing work, the robustness of the model can be further strengthened by concentrating on high confidence areas [29]. We introduce the High-confidence Nearest Pseudo Update (HNPU) Algorithm 1 to aggregate information from high-confidence neighborhoods for more robust adaptation. Specifically, for a given decoded point $\mathbf{x}$, we search the nearest $k$ neighbors in original coordinates $P$ with a constructed KD-tree $T$:
$\mathcal{N}_{k}(\mathbf{x},T)=\{\mathbf{x}^{\prime}\in P\mid\|\mathbf{x}-\mathbf{x}^{\prime}\|\leq d_{\text{th}},|\mathcal{N}_{k}|=k\},$ (12)
where $d_{\text{th}}$ is the distance threshold. Then the pseudo-label for $\mathbf{x}$ is updated with aggregated confidence scores:
$\hat{\mathbf{y}}(\mathbf{x})=\begin{cases}\mathbf{y}(\mathbf{x}),&C_{\text{conf}}(\mathbf{x})>\tau,\\ \mathbf{y}_{(}\mathbf{x}^{\prime}),&C_{\text{conf}}(\mathbf{x})\leq\tau\land\max\limits_{\mathbf{x}^{\prime}\in\mathcal{N}_{k}(\mathbf{x},P)}C_{\text{conf}}(\mathbf{x}^{\prime})>\tau,\\ -1,&\text{otherwise}.\end{cases}$ (13)
As shown in Fig. 4, a teacher-student network transfers information between the source and target domains during the cross-domain adaptation. This is done using the adaptation loss $\mathcal{L}_{tot}=\mathcal{L}_{s\to t}+\mathcal{L}_{t\to s}$, which captures consistent semantic information between the source and target domains. In the decoder branch, the adaptation process incorporates additional segmentation losses introduced by the decoded coordinates and the pseudo-labels updated by HNPU. The updated segmentation loss for the source domain is defined as :
$\mathcal{L}_{sd}=\mathcal{L}_{seg}\left(\Phi_{\theta}\left(\text{HNPU}(\mathbf{x}_{\mathcal{S}},\mathbf{y}_{\mathcal{S}})\right),L_{\mathcal{S}\text{u}}\right),$ (14)
and the updated segmentation loss for the target domain point clouds is denoted as:
$\mathcal{L}_{td}=\mathcal{L}_{seg}\left(\Phi_{\theta}\left(\text{HNPU}(\mathbf{x}_{\mathcal{T}},\hat{\mathbf{y}}_{\mathcal{T}})\right),L_{\mathcal{T}\mathrm{u}}\right),$ (15)
in which $L_{\mathcal{T}\text{u}}$ and $L_{\mathcal{S}\text{u}}$ are the pseudo-labels updated by HNPU with decoded coordinates and the confidence information for both domains. Intuitively, the decoder branch aligns the input distribution with the clean source distribution. The probability decoder transfers spatial information from the source domain, helping to mitigate adversarial shifts in input point clouds and reduce the domain gap for target domain data. In contrast, HNPU significantly mitigates the impact of anomalous source domain data, further enhancing the performance of cross-domain adaptation.
Algorithm 1 High-Confidence Nearest Pseudo Update (HNPU)
0:  Decoded coordinates $P_{\text{fine}}$, original coordinates $P$, original labels $L$, confidence scores $C_{\text{conf}}$, confidence threshold $\tau$, distance threshold $d_{\text{th}}$, number of neighbors $k$.
0:  Updated points $P_{\text{u}}$, updated labels $L_{\text{u}}$
1:  Construct a KD-tree $T=\text{KD-Tree}(P)$.
2:  for each point $\mathbf{x}\in P_{\text{fine}}$ do
3:     Find $k$ nearest neighbors $\mathcal{N}_{k}(\mathbf{x},T)$.
4:     Assign pseudo-label $\hat{y}(\mathbf{x})$ using Eq. (13).
5:  end for
6:  return  $P_{\text{u}}$, $L_{\text{u}}$
@@ V-A Experimental Settings
We perform adversarial cross-domain adaptation for 3D semantic segmentation tasks on SynLiDAR [3], SemanticPOSS [30], and SemanticKITTI [31] datasets. SynLiDAR is a large-scale synthetic dataset created using Unreal Engine comprising 198,396 annotated point clouds with 32 semantic classes. SemanticPOSS includes 2,988 real-world point clouds categorized into 14 semantic classes. SemanticKITTI is a comprehensive segmentation dataset derived from LiDAR acquisitions of the KITTI dataset, consisting of 43,552 annotated real-world point clouds with over 19 semantic classes. Following CosMix [32], we adopt the same training and validation splits. We select the pre-trained MinkUNet34 on a clean SynLiDAR dataset as our target semantic segmentation model. Specifically, hyper-parameters $\delta$ and $\epsilon$ used for generating adversarial examples are empirically set to 0.1 and 2, respectively. Hyper-parameters $\lambda_{d}$, $\lambda_{\mathcal{S}d}$, $\lambda_{\mathcal{T}d}$ are empirically setting to 0.1 and are fixed. We begin by training the probability decoder branch with 2048 sampled points. During cross-domain adaptation, we continue training the decoder branch with randomly sampled input point clouds to prevent over-fitting, with the learning rate empirically set to 0.0001. All experiments in this section are conducted using the PyTorch [33] library and are conducted on 4$\times$ NVIDIA GeForce RTX 4090.

### v3

@@ III The AdvSynLiDAR Dataset
Generally, point clouds in the source domain are susceptible to subtle perturbations and deformations, which can be caused by sensor noise, environmental variations, and preprocessing artifacts [23, 24]. This can significantly degrade model performance and increase distribution discrepancies across domains, leading to inferior 3D unsupervised cross-domain adaptation (UDA) performance for point cloud semantic segmentation. Despite the significance of this challenge, to the best of our knowledge, there is currently a lack of standardized methods for evaluating the impact of such perturbations during adaptation. To address these issues, we introduce the AdvSynLiDAR dataset, built upon the structured SynLiDAR dataset [3]. Specifically, our dataset focuses on understanding the impact of attacks on the source domain.
@@ III-A Imperceptible Perturbation Points
In real-world scenarios, adversarial perturbations are usually subtle and spatially dependent, making them difficult to be detected. To mimic realistic attacks, we propose a perceptually-aware perturbation strategy. This strategy applies more significant perturbations to distant regions while minimizing changes in dense, perceptually sensitive areas, enabling a more precise evaluation of UDA robustness under an adversarial source domain.
@@ III-B Noisy Semantic Labels
Manual annotations in 3D point clouds often contain label noise, especially for occluded, distant, or complex objects, where annotation errors occur more frequently [25]. While standard UDA approaches assume that source domain annotations are reliable, real-world scenarios often involve systematic annotation noise, where annotators or automated labeling systems confidently assign incorrect labels. To systematically evaluate the impact of such errors, we introduce high-confidence mislabeling into the source domain data. This approach generates the highest-confidence incorrect class as the adversarial target, mimicking real-world scenarios with noisy labels.
@@ IV-B1 Robust Long-Tail Loss (RLT Loss)
During the pre-training phase, in order to balance the advantages of KPS loss [5] in handling long-tail classes with the strengths of SoftDICELoss in managing head classes, we introduce the Robust Long-Tail loss (RLT loss). SoftDiceLoss is widely used in segmentation tasks to measure the overlap between predictions and ground truth labels. Initially designed for 2D classification tasks, KPS loss aims to enhance the separability of features by adjusting the boundaries between key points and non-key points in the feature space. To apply KPS loss to higher-dimensional 3D segmentation tasks, we make several improvements to the original KPS loss:
Geometric Importance in 3D Key Point Definition. We introduce geometric importance as the criteria for defining key points in 3D point clouds, allowing for better identification of critical points under attack conditions. This enhances the discriminative ability of KPS loss for long-tail classes.
@@
Dynamic Boundary Adjustment for Class Imbalance. In standard classification tasks, all classes share the same decision boundary, which may not be ideal for long-tail classes due to their sparse distribution and low sample count. However, 3D point clouds often exhibit severe class imbalance, especially for long-tail classes as depicted in Fig. 1. To address this issue, we introduce a dynamic boundary adjustment mechanism that adjusts each class’s boundary based on sample density and classification difficulty. In this way, the smaller the class size (i.e., long-tail classes), the larger the boundary adjustment term $m_{\mathbf{y}_{i}}$, giving these classes more decision space during classification.
In summary, the KPS loss in 3D tasks encourages more significant boundaries for key points, thereby promoting the separability of long-tail class features.
@@ IV-B2 Probability Decoder Branch
Probability Decoder. Inspired by VRCNet [6], we introduce a probability decoder to mitigate latent discrepancies in source inputs dynamically. As illustrated in Fig. 3, to simulate realistic noise, we generate perturbed inputs by adding Gaussian noise and random augmentations to the clean source coordinates.
The decoder branch reconstructs the clean point cloud in three stages: reconstruction, completion, and enhancement, which progressively restore global and local geometric details. Specifically, in the reconstruction stage, VRCNet encodes the input features into global features and the latent distribution. It then reconstructs a coarse point cloud using the decoding distribution.
Additionally, the parameter $\lambda$ serves as a weighting factor. During the completion process, the input point cloud is reconstructed using the global feature and the latent variable distribution derived from the perturbed and augmented input point cloud $\mathbf{x}_{g}$. During training, the posterior distribution is regularized to approach the prior distribution.
In the enhancement stage, a self-attention module and up-sampling layers are added to capture local features at different scales and restore the point cloud to finer granularity. This enables adaptive restoration of perturbed point clouds while preserving essential structural information, mitigating the risk of introducing additional noise or erroneous labels.
@@
High-Confidence Nearest Pseudo Update. Motivated by existing work, the robustness of the model can be further strengthened by concentrating on high confidence areas [27]. We introduce the High-confidence Nearest Pseudo Update (HNPU) Algorithm LABEL:alg:hnpu to aggregate information from high-confidence neighborhoods for more robust adaptation.
As shown in Fig. 4, a teacher-student network transfers information between the source and target domains during the cross-domain adaptation. This is done using the adaptation loss $\mathcal{L}_{tot}=\mathcal{L}_{s\to t}+\mathcal{L}_{t\to s}$, which captures consistent semantic information between the source and target domains. In the decoder branch, the adaptation process incorporates additional segmentation losses introduced by the decoded coordinates and the pseudo-labels updated by HNPU.
Intuitively, the decoder branch aligns the input distribution with the clean source distribution. The probability decoder transfers spatial information from the source domain, helping to mitigate adversarial shifts in input point clouds and reduce the domain gap for target domain data. In contrast, HNPU significantly mitigates the impact of anomalous source domain data, further enhancing the performance of cross-domain adaptation.
@@ V-A Experimental Settings
We perform adversarial cross-domain adaptation for 3D semantic segmentation tasks on SynLiDAR [3], SemanticPOSS [28], and SemanticKITTI [29] datasets. SynLiDAR is a large-scale synthetic dataset created using Unreal Engine comprising 198,396 annotated point clouds with 32 semantic classes. SemanticPOSS includes 2,988 real-world point clouds categorized into 14 semantic classes. SemanticKITTI is a comprehensive segmentation dataset derived from LiDAR acquisitions of the KITTI dataset, consisting of 43,552 annotated real-world point clouds with over 19 semantic classes. Following CosMix [30], we adopt the same training and validation splits.